The regular octagon is inscribed in a circle with a radius of 5 inches.Each side of the octagon has a length of 3 inches.
a. find the perimeter of the regular octagon
b. Draw apothem RT. Find the length of the apothem .
c. Find the area of the regular octagon.
Please help me to understand this question!!! Please and Thank You!

Answer by fractalier(2101)

If you draw the diagram and label the sides and the all radii, you can see that you have eight triangles. Notice that the apothem of each one is the height of each triangle. We can use the Pythagoren Theorem to find the apothem.
5^2 = h^2 + (3/2)^2
25 = h^2 + 9/4
h^2 = 91/4
h = sqrt(91) / 2
The perimeter is easy, just eight sides times 3 inches, or 24 inches.
The area is found by
A = (1/2)ap
A = (1/2)(sqrt(91) / 2)(24)
A = 6*sqrt(91)